Which equation is used in this problem,, Sample of 50 packages is selected , the mean amount is different from 8.17 ozs, then mean amount is 8.159 ounces and the standard deviation is 0.051 ozs. Is there evidence that the population meAN amount is different from 8.17 ozs by using .05 level of significance and determine the p-value and interpret its meaning.
Answer by stanbon(75887)

Which equation is used in this problem,,
-------------------
Sample of 50 packages is selected , the mean amount is different from 8.17 ozs, then mean amount is 8.159 ounces and the standard deviation is 0.051 ozs.
Is there evidence that the population meAN amount is different from 8.17 ozs by using .05 level of significance and determine the p-value and interpret its meaning.
----
Ho: u = 0.817
Ha: u is not equal to 0.817 (claim)
-----
x-bar = 8.159
t(8.159) = (8.159-8.17)/[0.051/sqrt(50)] = -1.5251
-----
p-value = 2*P(t < -1.5251 when df = 49) = 0.1337
----
Since the p-value is greater than 5%, fail to reject Ho.
The test results do not support the claim.
=======================
Cheers,
Stan H.
==============